OpenStudy (anonymous):

find the equivalent rate of compound interest if there is a 8% p.a simple interest on a principal that is $10,000 for 5 years. Is there a formula to find the equivalent rate of compounding interest?

OpenStudy (anonymous):

for finding the equivalent rate of compounding interest the amount in the two cases are to be equal..

OpenStudy (anonymous):

If the rate is 8% simple interest then what will be the rate in compound interest?

OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

assume the rate to be r%per annum and then proceed

OpenStudy (anonymous):

That's what I did

OpenStudy (anonymous):

??

OpenStudy (anonymous):

I used the compounding forma to find r

OpenStudy (anonymous):

I inserted $4000 which is simple interest into the equation

OpenStudy (anonymous):

10000*(1+r/100)^5 = 10000*(1+5*8/100) now find r

OpenStudy (anonymous):

you may use logarithm...

OpenStudy (anonymous):

Ill solve it on paper and get back to you

OpenStudy (anonymous):

Where did you get 5 x 8

OpenStudy (anonymous):

from the simple interest formula for 5yrs @8%

OpenStudy (anonymous):

Yes but what formula did you use

OpenStudy (anonymous):

Why isn't the 5 represented as n

OpenStudy (anonymous):

I got 7%

OpenStudy (anonymous):

P(1+R)^n = P + P * r * t

OpenStudy (anonymous):

Is this correct?

OpenStudy (yttrium):

Just take note that F = P (1+i)^n Where P = principal value, F = future value, i = interest rate.

OpenStudy (anonymous):

P(1+R)^n = P + P * r * t is correct...

OpenStudy (anonymous):

Thanks @matricked